Ive been getting the wrong answers when calcuating these. Please help

21.50 mL of 0.320 M NaOH was added to 20.00 mL of
Answer #1

4)

Molarity of mixer = (20*0.41+21.5*0.32)/41.5

   = 0.36 M

in the mixture , HCl is excess.

so that

pH = -log(0.36) = 0.44

[H+] = 0.36 M

[OH-] = 10^(-14)/0.36 = 2.78*10^-14 M


5)

pH of HCl = -log0.001 = 3

as the pH is less than 4. we sholud not pour the things down the drain